The time it takes to travel a fixed distance varies inversely as the rate of travel. If a car can cover a certain distance in 7 hours at 40 miles per hour, how fast would one need to travel in order to cover the distance in 4 hours?
Answer by Cromlix(4381)

Time varies inversely as rate of travel
T = k/R
(k is the constant of variation)
Use given example to find k
7 = k/40
Cross multiply
280 = k
So, formula = T = 280/R
Using formula:
T = 280/R
4 = 280/R
R = 280/4
R = 70 mph.
Hope this helps.
:-)